2. Solve the following systems of equations by elimination. Verify the solutions.
a. r + 2s + 1 = 0
r + 5s + 28 = 0

b. 4m - 3n = 27
8m - 6n = 18



c. 0.6x = 1.2 + 0.3y
2.2x - 1.8y - 1.6 = 0

3. Frank has $8 000 that he plans to split into two investments. He wrote the following two equations to represent the interest he will earn from each of the two investment options.
• 2000A + 6000B = 520
• 4000A + 4000B = 480
Determine the interest rates, A and B, as percentages.

Answer by FrankM(1040)


r + 5s + 28 = 0
r + 2s + 1 = 0 subtract from first equation.
3s+27=0
s=-9
r + 2(-9) + 1 = 0
r -18+1 =0
r=17